The age of the population in Wakanda is shown in the following table in the form of a grouped frequency distribution.

  Age (years)

  Number (‘000)

 0 Under 5

5 under 15 15 under 30 30 under 50 50 under 65 65 under 80 80 under 100

Total

Use the data in the table above to estimate for of population in Wakanda the:

1 375.3 2 750.5 4 514.7 6 137.7 3 821.7 2 045.7 783.2

        100 under 110

what is the median age?


1
Expert's answer
2021-05-11T14:34:35-0400

Table can be formed as:




Median number (m) = n/2= 21431.9/2

m = 10715.95


So, median class = 30 - 50


Median age "M=l_1+\dfrac{(n/2-cf)}{f}\cdot l"


where "l_1" = lower number of median class

n/2=median number

cf = cumulative frequency of class preceding to median class

f = frequency/number of median class

i = class length


and we have

"l_1" = 30

n/2 = 10715.95

cf = 8640.5

f = 6137.7

i = 20


Therefore, Median "M=l_1+\dfrac{(n/2-cf)}{f}\cdot l"

"=30+[\dfrac{10715.95-8640.5}{6137.7}]\cdot 20\\=30+6.76\\=36.76"


Hence, median age = 36.76 years